For the first order linear ODE:

y'(x)+p(x)y(x)=r(x) and let D(y(x))=y'(x)+p(x)y(x),

Find the solution of the following ODE:

y'+y(x^5)=cos(wx), y(0)=0
Answer #1

y + y(+5) = cos(wa) Finst onder linear o DE 10) = 0 i plas= 25, Qing= cos(oo) Integrating factor socdx yx (T. F) = ( 8m x (II + 36w? I = 6 cos (wa) et + 36w sin order I (14 36w) - 6 costwn) e + 36 w sincwadie I = (6 coswn + 36 w Sincwad je z 1+36w?